OpenStudy (anonymous):

Find the value(s) of guaranteed by the Mean Value Theorem for Integrals for the function over the given interval. @NCount

OpenStudy (anonymous):

\[f(x)=\frac{ 9 }{ x ^{3} }\] [1,3]

OpenStudy (anonymous):

\[f(c) = \frac{1}{3-1} \int\limits_{1}^{3} \frac{9}{x^2} dx\] solve for c

OpenStudy (anonymous):

Thanks dude!

OpenStudy (anonymous):

this is the mean value theorem for INTEGRAL

OpenStudy (anonymous):

Whoops! Thanks sourwing. I saw MVT and started running without reading that bit.

OpenStudy (anonymous):

so how did you get the \[\frac{ 1 }{ 1-3 }\]

OpenStudy (anonymous):

Here is the general formula sourwing pulled that from: \[f(c)=\frac{ 1 }{ b-a }\int\limits_{b}^{a}f(x) dx\]

OpenStudy (anonymous):

EDIT: the integral should be from a to b, NOT b to a.

OpenStudy (anonymous):

Thank you thank you

OpenStudy (anonymous):

Sure thing!